# Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy

### v3

@@ 1 Introduction
We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is associated with a theme (the set of themes derived from previous studies (Islam and Goldwasser, 2024; Islam et al., 2023b; Pacheco et al., 2023; Islam and Goldwasser, 2022; Pacheco et al., 2022a)). An overview of our framework is illustrated in Figure. 1. Our framework provides a way of clustering talking points motivated by themes. We have assigned instances with themes. For each theme cluster, we cluster the associated instances using a clustering algorithm. We characterize the sub-clusters by performing zero-shot multi-document summarization of the top-K instances assigned to each sub-cluster. Then, we prompt LLMs in a zero-shot manner to generate the talking point advocating in the sub-cluster summary in the context of the given theme. For redundancy check, we assess the similarity between talking point pairs and merge overlapping talking points based on threshold. We have an optional human evaluation phase for newly generated talking points. Finally, the assignment phase maps instances to their most likely talking points.
@@ 3.5 Mapping Instances to Arguments
Table 1: Coverage and mapping quality w.r.t. Human Judgements.
@@ 4 Case Studies
We explore two case studies involving discussions on social media: (1) climate campaigns and (2) COVID-19 vaccine campaigns. For climate campaigns, we work on the corpus of $14k$ ads released by Islam et al. (2023b). The dataset focuses on climate-related English ads on Facebook based on the US from January $2021$ to January $2022$. All ads in this corpus contain predictions for stance (e.g., pro-energy, clean-energy) and theme (e.g., support climate policy.) (Islam and Goldwasser, 2024). For COVID-19 vaccine campaigns, we use the corpus of $9k$ ads released by Islam and Goldwasser (2022) focusing on COVID-19 vaccine-related English ads on Facebook based on US from December $2020$ to January $2022$. All ads in this corpus contain predictions for moral foundation (e.g., care/harm) (Haidt and Graham, 2007) and theme (e.g., vaccine equity.) (Islam and Goldwasser, 2022). For each ad of both corpora, there are the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, distribution over impressions broken down by gender (male, female, unknown), age (7 groups), and location down to states in the USA. Derived themes are shown in Table 5 in Appendix A.1. Data statistics regarding the number of instances in each theme for climate (Fig. 4(a)) and COVID-19 vaccine (Fig. 4(b)) campaign are shown in Fig. 4 in App. A.2. Additional details about the dataset can be found in the original publication.
@@
Our main goal is to use our LLMs-in-the-Loop framework to identify prominent arguments in the corpora introduced above. At the beginning, we have ads and their associated themes. Those themes are derived by Islam and Goldwasser (2024, 2022). At the first iteration, we follow the step introduced in Section 3 to obtain the list of talking points under each theme and the mapping $ads\rightarrow arguments$. Hyperparameters are detailed in App. A.3. After iteration $1$, we identify $113$ arguments for Climate campaigns and $47$ arguments for COVID-19 vaccine campaigns (Table 1). In $2^{nd}$ iteration, we repeat this process for the rest of the ads remaining unassigned from iteration $1$ (by applying a threshold of $<0.5$), resulting in the discovery of $100$ and $31$ additional arguments for climate and COVID-19 vaccine campaigns respectively (Table 1). An example of a resulting set of arguments under the ‘Patriotism’ (one of the themes from climate campaigns) and ‘VaccineEquity’ (one of the themes from COVID-19 vaccine campaigns) themes after two rounds of iterations of our LLMs-in-the-Loop are shown in Table 8 in App. A.4. We show the number of sub-clusters identified under each theme for both iterations of the climate (Table 6) and COVID-19 (Table 7) case studies are provided in App. A.2. The prompt templates are provided in Figure 5 in App. A.5. A concrete example of prompt for summarizing top-5 instances under patriotism theme of climate campaign and the generated summary are shown in Fig. 6 in App. A.5. Prompt example of generating talking point from a summary of top-5 instances under patriotism theme of climate campaign dataset is shown in Fig. 7 in App. A.5. The complete set of final arguments is provided in the App. A.6 in Table LABEL:arg_climate and Table LABEL:arg_covid.
@@ 4.1 Evaluation
To evaluate the performance of our ad-to-argument mapping, we sort the ads according to their semantic distance from their assigned arguments. We then compute the three quartiles and sample a set of $12$ ads per theme, such that $3$ ads are randomly sampled from each quartile. Subsequently, we manually annotate whether the mapping is correct or not (detailed in Appendix A.7). This process results in $300$ ads in the $1^{st}$ iteration and another $300$ ads from the $2^{nd}$ iteration from the climate case study. On the other hand, from the COVID-19 vaccine case study, we obtain $168$ ads in the $1^{st}$ iteration and another $168$ ads from the $2^{nd}$ iteration.
@@ 4.2 Argumentative Cohesion Comparison
For the climate debate case study, we perform a correlation test between the identified talking points and the stance expressed in the ads (i.e., pro-energy or clean-energy). We calculate the Pearson correlation matrices and present them in Fig. 2. We compare our discovered talking points with a set of topics extracted using LDA. In Fig. 2(f), we show random $25$ arguments (among $213$ total arguments of climate) to show the correlation between the arguments and the stance. To compare with LDA baseline, we choose $10$ topics (Fig. 2(d)) and $30$ topics (Fig. 2(e)) to keep our $25$ arguments as middle ground. We observe that our talking points (Fig. 2(f)) have stronger correlations with stances than the derived LDA topics (Fig. 2(d), Fig. 2(e)). For example, we note that pro-energy stance is strongly correlated with talking point such as “Criticizes President Biden’s economic policies for significantly increasing gas prices and living costs, urging a shift in leadership or policy.”
Hyperparameters are detailed in App. A.3.
@@ 6.1 Demographic Targeting
Understanding how messages are tailored to target particular demographic groups is crucial for opinion analysis. To analyze the variation in talking points on climate campaigns across different age groups, we examine targeted ads within three age categories, i.e., young people (ages $13$-$24$), working-age people (ages $25$-$54$), and the older population (age $55+$), focusing on California (CA) and Texas (TX). We choose these states because TX is recognized as the world’s energy capital333www.eia.gov//todayinenergy/detail.php?id=49356, while CA has recently positioned itself as a leading advocate in the battle against climate change444www.pewtrusts.org//en/research-and-analysis/blogs/stateline/2022/10/06/california-takes-leading-edge-on-climate-laws-others-could-follow. To identify the top-k most frequently mentioned entities in the texts, we prompt LLMs in a zero-shot setting. We show the results in App. A.8 (Table 9). We observe that talking points vary by age group in TX. For young people, the focus is on energy reliability and affordability, learning from Texas’ deregulation disaster. For the working-age population, the emphasis is on accelerating the transition to renewable energy to combat climate change and support economic growth through job creation. Meanwhile, for the older population, the discussion centers on the need for urgent legislative action on climate change.
@@ 6.2 Shifts in Talking Points Following Significant Events
Federal Vaccine Mandate: On September $09$, $2021$, President Biden announced an executive order on COVID-19 vaccine mandates for federal workers, large employers, and health care staff888www.whitehouse.gov/briefing-room/presidential-actions/2021/09/09/executive-order-on-requiring-coronavirus-disease-2019-vaccination-for-federal-employees/. Figure 3(a) displays the top-four shifts in talking points in sponsored advertisements over the five months before and after Hurricane Ida struck the USA. There is a noticeable presence of arguments for "urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future." as well as advocacy for "embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility." following the hurricane (see Figure 3(a)). Figure 3(b) illustrates the talking points used in sponsored ads three days before and after President Biden announced the federal vaccine mandate. In the aftermath, three new predominant arguments emerged: "Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.", "COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.", "Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support." (Figure 3(b)). These shifts suggest that sponsored content is highly responsive to real-world events, quickly reflecting and potentially shaping public discourse on pivotal issues.
@@ A.2 Data Statistics
Table 8: Example of resulting arguments after $1^{st}$ and $2^{nd}$ iteration. For climate campaigns, arguments are shown under the ‘Patriotism’ theme. For COVID-19 vaccine campaigns, arguments are shown under the ‘VaccineEquity’ theme. All themes derived by previous studies are shown in Appendix A.1 (Table 5)
@@ A.4 Arguments under specific Themes
Figure 5: Prompt templates (shown as zero-shot).
A.5 Prompting
In this section, we show the prompt templates used in our work to generate a sub-cluster summary (Figure. 5 (a)) and a talking point from each summarized sub-cluster (Figure. 5 (b)). Prompt examples for summarizing top-5 instances and generating talking point from that summary under patriotism theme of climate campaign dataset are shown in Fig. 6 and Fig. 7 respectively.
A.6 Resulting Arguments
@@ A.6.1 Climate Campaigns
A.6.2 COVID-19 Vaccine Campaigns

### v4

@@ 1 Introduction
We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is associated with a theme (the set of themes derived from previous studies (Islam and Goldwasser, 2024; Islam et al., 2023b; Pacheco et al., 2023; Islam and Goldwasser, 2022; Pacheco et al., 2022a)). An overview of our framework is illustrated in Fig. 1. Our framework provides a way of clustering talking points motivated by themes. We have assigned instances with themes. For each theme cluster, we cluster the associated instances using a clustering algorithm. We characterize the sub-clusters by performing zero-shot multi-document summarization of the top-K instances assigned to each sub-cluster. Then, we prompt LLMs in a zero-shot manner to generate the talking point advocating in the sub-cluster summary in the context of the given theme. For redundancy check, we assess the similarity between talking point pairs and merge overlapping talking points based on threshold. We have an optional human evaluation phase for newly generated talking points. Finally, the assignment phase maps instances to their most likely talking points.
@@ 3.5 Mapping Instances to Arguments
Table 1: Coverage and mapping quality w.r.t. Human Judgments.
@@ 4 Case Studies
We explore two case studies involving discussions on social media: (1) climate campaigns and (2) COVID-19 vaccine campaigns. For climate campaigns, we work on the corpus of $14k$ ads released by Islam et al. (2023b). The dataset focuses on climate-related English ads on Facebook based on the US from January $2021$ to January $2022$. All ads in this corpus contain predictions for stance (e.g., pro-energy, clean-energy) and theme (e.g., support climate policy.) (Islam and Goldwasser, 2024). For COVID-19 vaccine campaigns, we use the corpus of $9k$ ads released by Islam and Goldwasser (2022) focusing on COVID-19 vaccine-related English ads on Facebook based on US from December $2020$ to January $2022$. All ads in this corpus contain predictions for moral foundation (e.g., care/harm) (Haidt and Graham, 2007) and theme (e.g., vaccine equity.) (Islam and Goldwasser, 2022). For each ad of both corpora, there are the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, distribution over impressions broken down by gender (male, female, unknown), age (7 groups), and location down to states in the USA. Derived themes are shown in Table 5 in App. A.1. Data statistics regarding the number of instances in each theme for climate (Fig. 7(a)) and COVID-19 vaccine (Fig. 7(b)) campaign are shown in Fig. 7 in App. A.2. Additional details about the dataset can be found in the original publication.
@@
Our main goal is to use our LLMs-in-the-Loop framework to identify prominent arguments in the corpora described earlier. At the beginning, we have ads and their associated themes. Those themes are derived by Islam and Goldwasser (2024, 2022). At the first iteration, we follow the step introduced in Section 3 to obtain the list of talking points under each theme and the mapping $ads\rightarrow arguments$. Hyperparameters are detailed in App. A.3. After iteration $1$, we identify $113$ arguments for Climate campaigns and $47$ arguments for COVID-19 vaccine campaigns (Table 1). In $2^{nd}$ iteration, we repeat this process for the rest of the ads remaining unassigned from iteration $1$ (by applying a threshold of $<0.5$), resulting in the discovery of $100$ and $31$ additional arguments for climate and COVID-19 vaccine campaigns respectively (Table 1). An example of a resulting set of arguments under the ‘Patriotism’ (one of the themes from climate campaigns) and ‘VaccineEquity’ (one of the themes from COVID-19 vaccine campaigns) themes after two rounds of iterations of our LLMs-in-the-Loop are shown in Table 8 in App. A.4. We show the number of sub-clusters identified under each theme for both iterations of the climate (Table 6) and COVID-19 (Table 7) case studies are provided in App. A.2. The complete set of final arguments is provided in the App. A.5 in Table LABEL:arg_climate and Table LABEL:arg_covid.
Figure 2: Prompt templates (shown as zero-shot).
4.1 Prompting
In this section, we show the prompt templates used in our work to generate a sub-cluster summary (Fig. 2 (a)) and a talking point from each summarized sub-cluster (Fig. 2 (b)). A concrete example of a prompt for summarizing top-5 instances under patriotism theme of climate campaign and the generated summary are shown in Fig. 3. A prompt example of generating a talking point from a summary of top-5 instances under patriotism theme of the climate campaign dataset is illustrated in Fig. 4.
@@ 4.2 Evaluation
To evaluate the performance of our ad-to-argument mapping, we sort the ads according to their semantic distance from their assigned arguments. We then compute the three quartiles and sample a set of $12$ ads per theme, such that $3$ ads are randomly sampled from each quartile. Subsequently, we manually annotate whether the mapping is correct or not (detailed in App. A.6). This process results in $300$ ads in the $1^{st}$ iteration and another $300$ ads from the $2^{nd}$ iteration from the climate case study. On the other hand, from the COVID-19 vaccine case study, we obtain $168$ ads in the $1^{st}$ iteration and another $168$ ads from the $2^{nd}$ iteration.
@@ 4.3 Argumentative Cohesion Comparison
For the climate debate case study, we perform a correlation test between the identified talking points and the stance expressed in the ads (i.e., pro-energy or clean-energy). We calculate the Pearson correlation matrices and present them in Fig. 5. We compare our discovered talking points with a set of topics extracted using LDA. In Fig. 5(f), we show random $25$ arguments (among $213$ total arguments of climate) to show the correlation between the arguments and the stance. To compare with LDA baseline, we choose $10$ topics (Fig. 5(d)) and $30$ topics (Fig. 5(e)) to keep our $25$ arguments as middle ground. We observe that our talking points (Fig. 5(f)) have stronger correlations with stances than the derived LDA topics (Fig. 5(d), Fig. 5(e)). For example, we note that pro-energy stance is strongly correlated with talking point such as “Criticizes President Biden’s economic policies for significantly increasing gas prices and living costs, urging a shift in leadership or policy.” Hyperparameters are detailed in App. A.3.
@@ 6.1 Demographic Targeting
Understanding how messages are tailored to target particular demographic groups is crucial for opinion analysis. To analyze the variation in talking points on climate campaigns across different age groups, we examine targeted ads within three age categories, i.e., young people (ages $13$-$24$), working-age people (ages $25$-$54$), and the older population (age $55+$), focusing on California (CA) and Texas (TX). We choose these states because TX is recognized as the world’s energy capital333www.eia.gov//todayinenergy/detail.php?id=49356, while CA has recently positioned itself as a leading advocate in the battle against climate change444www.pewtrusts.org//en/research-and-analysis/blogs/stateline/2022/10/06/california-takes-leading-edge-on-climate-laws-others-could-follow. To identify the top-k most frequently mentioned entities in the texts, we prompt LLMs in a zero-shot setting. We show the results in App. A.7 (Table 9). We observe that talking points vary by age group in TX. For young people, the focus is on energy reliability and affordability, learning from Texas’ deregulation disaster. For the working-age population, the emphasis is on accelerating the transition to renewable energy to combat climate change and support economic growth through job creation. Meanwhile, for the older population, the discussion centers on the need for urgent legislative action on climate change.
@@ 6.2 Shifts in Talking Points Following Significant Events
Federal Vaccine Mandate: On September $09$, $2021$, President Biden announced an executive order on COVID-19 vaccine mandates for federal workers, large employers, and health care staff888www.whitehouse.gov/briefing-room/presidential-actions/2021/09/09/executive-order-on-requiring-coronavirus-vaccination-for-federal-employees/. Figure 6(a) displays the top-four shifts in talking points in sponsored advertisements over the five months before and after Hurricane Ida struck the USA. There is a noticeable presence of arguments for "urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future." as well as advocacy for "embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility." following the hurricane (see Figure 6(a)). Figure 6(b) illustrates the talking points used in sponsored ads three days before and after President Biden announced the federal vaccine mandate. In the aftermath, three new predominant arguments emerged: "Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.", "COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.", "Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support." (Figure 6(b)). These shifts suggest that sponsored content is highly responsive to real-world events, quickly reflecting and potentially shaping public discourse on pivotal issues.
@@ 9 Ethics Statement
10 Acknowledgments
We thank the anonymous reviewers for their insightful comments. This work was partially funded by Purdue Graduate School Summer Research Grant (to TI) and NSF CAREER award IIS-2048001.
@@ A.2 Data Statistics
Table 8: Example of resulting arguments after $1^{st}$ and $2^{nd}$ iteration. For climate campaigns, arguments are shown under the ‘Patriotism’ theme. For COVID-19 vaccine campaigns, arguments are shown under the ‘VaccineEquity’ theme. All themes derived by previous studies are shown in App. A.1 (Table 5)
@@ A.4 Arguments under specific Themes
A.5 Resulting Arguments
@@ A.5.1 Climate Campaigns
A.5.2 COVID-19 Vaccine Campaigns